A train moving west with an initial velocity of 20 m/s accelerates at 4 m/s2 for 10 seconds. During this time, the train moves a distance of
meters.
1
Expert's answer
2020-09-06T17:17:11-0400

The distance is


"x=v_0t+\frac{at^2}{2},\\\space\\
x=20\cdot10+\frac{4\cdot10^2}{2}=400\text{ m}.\\\space\\"